OpenStudy (anonymous):

The volume of V of cube with edge length "s" is givem by the function V(s) = s^3. Find V(4). Explain what V(4) represents.

OpenStudy (dan815):

if V(s)=s^3 then V(4)=4^3

OpenStudy (dan815):

It was a function of s, and s happened to be 4 so wherever you see an s you replace that with 4

OpenStudy (dan815):

4^3=64